### v1

@@ 2 Related Work
Hybrid retrieval frameworks can combine sparse and dense retrieval by adding a subsequent re-ranking stage to merge their results and improve retrieval quality. Neural re-rankers [23] have demonstrated substantial improvements in ranking accuracy across multiple domains. Recently, large language models (LLMs) have been employed as cross-encoders, jointly encoding claim–document pairs to capture nuanced semantic relationships [11, 12]. In this work, we adopt such a hybrid retrieval architecture by combining sparse retrieval via BM25, dense neural retrieval, and LLM-based re-ranking.
@@ 3.1 Lexical Retrieval
In contrast to the baseline BM25 provided by the challenge organizers [5], we apply additional normalization steps to improve match quality. Our pipeline includes lowercasing, punctuation removal, and subword tokenization using byte pair encoding (BPE) [24]. We chose subword tokenization over lemmatization to maximize n-gram overlaps between informal query terms and formal document vocabulary. Hashtags are removed, while symbols such as percentages (%) are preserved to maintain scientific meaning. We detail additional LLM-based pre-processing experiments, which were excluded from the pipeline, in Appendix A.
@@
At inference time, the BM25 retriever returns the top-$30$ documents ranked by relevance. This candidate set provides strong lexical matches for downstream re-ranking and complements the semantic retriever. Empirically, we found that increasing the number of retrieved candidates beyond $30$ did not lead to improved retrieval performance, but substantially increased computational cost during re-ranking.
@@ 3.2 Semantic Retrieval
We initialize our dense retriever with the INF-Retriever-v1 model [10], a fine-tuned variant of gte-Qwen2-7B-instruct [26], and further fine-tune it on the CLEF CheckThat! training set using the multiple negatives ranking (MNR) loss [27]. The model is trained to assign higher similarity scores to (social media post, document) pairs with known associations than to randomly sampled negatives.
Fine-tuning is performed with a maximum input length of $8,192$ tokens to accommodate long documents without truncation. Inputs (queries and documents) are tokenized independently, and embeddings are computed via last-token pooling. We apply LoRA adapters [28] to the final eight transformer layers to reduce memory and training time [29] and optimize using AdamW [30] with cosine learning rate decay and gradient accumulation. Full details of the fine-tuning setup, including model initialization, LoRA configuration, optimization strategy, and evaluation protocol, are provided in Appendix B.
@@
We precompute document embeddings and store them in a FAISS index [8, 9]. The embeddings are normalized using the L2 norm, allowing cosine similarity to be computed efficiently via dot products. At inference time, the social media post is encoded into a dense vector using the same model, and the top $100$ most similar documents are retrieved. We avoid chunking abstracts, as empirical results have shown that full-document retrieval performs better.
@@ 3.3 Re-Ranking
While dense and sparse retrievers are computationally faster, the subsequent re-ranking process is computationally intensive. Unlike embedding models, which independently embed each document and query into vectors and compute similarity using a distance metric, the re-ranker processes query-document pairs to directly output a similarity score. The computational cost of these pairwise comparisons limits re-ranking to small candidate subsets, making the initial retrieval stage essential for filtering documents.
@@
We conducted experiments with various re-ranking models (see Appendix C for discussion). Based on these experiments, we select BAAI/bge-reranker-v2-gemma [11, 12], an LLM-based cross encoder built on Gemma [31]. To balance cost and performance, we re-rank a candidate pool consisting of the top $100$ documents from dense retrieval and the top $30$ from BM25 retrieval. We favor the former due to its stronger individual performance (see Section 4). This set of candidates is scored by the re-ranker, and the top five results are returned as the pipeline output.
@@ 4 Results
On the development set, employing INF-Retriever-v1 yields an absolute improvement of $10.03$ percentage points in MRR@5 over the BM25 baseline. Fine-tuning the retriever further increases MRR@5 by $1.98$ points, reaching a final score of $67.19$%. In terms of Precision@100, the base model achieves a $13.7$-point gain compared to the BM25 baseline, with fine-tuning contributing an additional $1.1$-point improvement. These gains are similar on the test set: INF-Retriever-v1 improves MRR@5 by $11.4$ points over the BM25 baseline, and fine-tuning adds a further $2.2$-point gain, culminating in an MRR@5 of $56.72$%. Precision@100 follows a similar trend, with respective gains of $17.85$ and $1.93$ percentage points. These consistent improvements across both development and test splits highlight the effectiveness and robustness of semantic retrieval, particularly when fine-tuning is applied.
We also experimented with data augmentation techniques, including HyDE-generated queries and alternative document variants. However, these did not yield further gains. A discussion on data augmentation is provided in Appendix D.
@@ 5 Discussion and Conclusion
    In-domain fine-tuning yields measurable gains: Fine-tuning the dense retriever improved MRR@5 by approx. $+2$ percentage points and led to a Precision@100 of $89.21$%. While pre-trained models perform well out of the box, domain adaptation further improves alignment between informal queries and scientific abstracts.

### v2

@@ 2 Related Work
Hybrid retrieval frameworks can combine sparse and dense retrieval by adding a subsequent re-ranking stage to merge their results and improve retrieval quality. Neural re-rankers [23] have demonstrated substantial improvements in ranking accuracy across multiple domains. Recently, large language models (LLMs) have been employed as cross-encoders, jointly encoding claim–document pairs to capture nuanced semantic relationships [11, 12]. In this work, we adopt such a hybrid retrieval architecture by combining sparse retrieval via BM25, dense neural retrieval, and LLM-based re-ranking to leverage the different strengths of these retrieval methods.
@@ 3.1 Lexical Retrieval
In contrast to the baseline BM25 provided by the challenge organizers [5], we apply additional normalization steps to improve match quality. Our pipeline includes lowercasing, punctuation removal, and subword tokenization using byte pair encoding (BPE) [24]. We chose subword tokenization over lemmatization to maximize n-gram overlaps between informal query terms and formal document vocabulary. Hashtags are removed, while symbols such as percentages (%) are preserved to maintain scientific meaning. This design choice specifically addresses the domain gap between informal social media queries and formal scientific language by increasing the chance of partial term matches. We detail additional pre-processing experiments, which were excluded from the pipeline, in Appendix A.
@@
At inference time, the BM25 retriever returns the top-$30$ documents ranked by relevance. This candidate set provides strong lexical matches for downstream re-ranking and complements the semantic retriever.
@@ 3.2 Semantic Retrieval
We initialize our dense retriever with the INF-Retriever-v1 model [10], a fine-tuned variant of gte-Qwen2-7B-instruct [26], chosen for its strong long-text retrieval performance and open-source availability. We fine-tune it on the CLEF CheckThat! training set using the multiple negatives ranking (MNR) loss [27], training it to assign higher similarity scores to (social media post, document) pairs with known associations than to random negatives.
Fine-tuning uses a maximum input length of $8,192$ tokens to avoid truncation. Queries and documents are tokenized independently, embeddings use last-token pooling, and LoRA adapters [28] are applied to the final eight transformer layers to reduce memory and training time [29]. We optimize with AdamW [30] using cosine learning rate decay and gradient accumulation. Full details of the fine-tuning setup are provided in Appendix B.
@@
We pre-compute document embeddings and store them in a FAISS index [8, 9]. The embeddings are normalized using the L2 norm, allowing cosine similarity to be computed efficiently via dot products. At inference time, the social media post is encoded into a dense vector using the same model, and the top $100$ most similar documents are retrieved. We avoid chunking abstracts, as empirical results have shown that full-document retrieval performs better.
@@ 3.3 Re-Ranking
While dense and sparse retrievers are computationally faster, the subsequent re-ranking process is computationally intensive. Unlike embedding models, which independently embed each document and query into vectors and compute similarity using a distance metric, the re-ranker processes each query-document pair jointly to directly output a similarity score. The computational cost of these pairwise comparisons limits re-ranking to small candidate subsets, making the initial retrieval stage essential for filtering documents.
@@
We evaluated various re-ranking models (see Appendix C) and selected BAAI/bge-reranker-v2-gemma [11, 12], an LLM-based cross encoder built on Gemma [31], as it performed best. To balance cost and performance, we re-rank the top $100$ dense and top $30$ BM25 candidates, favoring the former due to stronger individual performance (see Section 4). Empirically, increasing the number of BM25 candidates beyond $30$ did not improve re-ranking performance but substantially increased computational cost, whereas increasing the dense candidates to $100$ led to substantial gains. After removing duplicates within the $130$ candidates for each query, the re-ranker scores all remaining candidates from scratch, and the top five results are returned as output.
@@ 4 Results
On the development set, employing INF-Retriever-v1 yields an absolute improvement of $10.03$ percentage points in MRR@5 over the BM25 baseline. Fine-tuning the retriever further increases MRR@5 by $1.98$ points, reaching a final score of $67.19$%. In terms of Precision@100, the base model achieves a $13.7$-point gain compared to the BM25 baseline, with fine-tuning contributing an additional $1.1$-point improvement. These gains are similar on the test set: INF-Retriever-v1 improves MRR@5 by $11.4$ points over the BM25 baseline, and fine-tuning adds a further $2.2$-point gain, culminating in an MRR@5 of $56.72$%. Precision@100 follows a similar trend, with respective gains of $17.85$ and $1.93$ percentage points. These consistent improvements across both development and test splits highlight the effectiveness and robustness of semantic retrieval, particularly when fine-tuning is applied. We also experimented with data augmentation techniques, including HyDE-generated queries and alternative document variants. However, these did not yield further gains. A discussion on data augmentation is provided in Appendix D.
@@ 5 Discussion and Conclusion
    In-domain fine-tuning improves performance: Fine-tuning the dense retriever improved MRR@5 by approx. $+2$ percentage points and led to a Precision@100 of $89.21$%. While pre-trained models perform well out of the box, domain adaptation further improves alignment between informal queries and scientific abstracts.